OpenStudy (anonymous):

Factor the trinomial completely. 12x^2-13x-14

OpenStudy (anonymous):

so to factor this one i looked at the 12 and thought of ways i could break that up 12, 1 6,2 3,4 I started with the 3,4 (3x+?)(4x+?) This way happened to be the one that worked I got a final answer of (3x+2)(4x-7)

OpenStudy (anonymous):

(4x-7)(3x+2)

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

(3x+2)(4x-7) 3x * 4x = 12x^2 For the first part = 12x^2 2 * 4x = 8x 3x * -7 = -21x For the second part -21x + 8x = -13x 2 * -7 = -14 For the third part = -14 Giving you 12x^2-13x-14